OpenStudy (anonymous):

HELP PLEASE!!! WILL FAN & GIVE MEDAL!! Solve the equation. |k + 6| = 3 A) {-3, 9} B.) {-3, 3} c) {-9, -3} d.) {all real numbers greater than or equal to –9 and less than or equal to –3}

OpenStudy (anonymous):

@beccaboo333 @Xmoses1

OpenStudy (texaschic101):

| k + 6 | = 3 k + 6 = 3 k = 3 - 6 k = -3 -(k + 6) = 3 -k - 6 = 3 -k = 3 + 6 -k = 9 k = -9 solution : (-9,-3)

OpenStudy (anonymous):

thanks!!